A bank loaned out $22,500, part of it at the rate of 3% annual interest, and the rest at 11% annual interest. The total interest earned for both loans was $1,915.00. How much was loaned at each rate?
Answer by ankor@dixie-net.com(22740)

A bank loaned out $22,500, part of it at the rate of 3% annual interest,
and the rest at 11% annual interest.
The total interest earned for both loans was $1,915.00.
How much was loaned at each rate?
:
let x = amt loaned at 11%
The total was 22500, therefore
(22500-x) = amt loaned at 3%
:
The interest equation
.11x + .03(22500-x) = 1915
distribute .03
.11x + 675- .03x = 1915
.11x - .03x = 1915 - 675
.08x = 1240
x = 1240/.08
x = $15,500 loaned at 11%
and
22500-15500 = $7,000 loaned at 3%
:
:
Check this
.11(15500) = 1705
.03(7000) = 210
-------------------
total int: 1915